# Detect Anything 3D in the Wild

### v1

@@ Abstract
Despite the success of deep learning in close-set 3D object detection, existing approaches struggle with zero-shot generalization to novel objects and camera configurations. We introduce DetAny3D, a promptable 3D detection foundation model capable of detecting any novel object under arbitrary camera configurations using only monocular inputs. Training a foundation model for 3D detection is fundamentally constrained by the limited availability of annotated 3D data, which motivates DetAny3D to leverage the rich prior knowledge embedded in extensively pre-trained 2D foundation models to compensate for this scarcity. To effectively transfer 2D knowledge to 3D, DetAny3D incorporates two core modules: the 2D Aggregator, which aligns features from different 2D foundation models, and the 3D Interpreter with Zero-Embedding Mapping, which mitigates catastrophic forgetting in 2D-to-3D knowledge transfer. Experimental results validate the strong generalization of our DetAny3D, which not only achieves state-of-the-art performance on unseen categories and novel camera configurations, but also surpasses most competitors on in-domain data. DetAny3D sheds light on the potential of the 3D foundation model for diverse applications in real-world scenarios, e.g., rare object detection in autonomous driving, and demonstrates promise for further exploration of 3D-centric tasks in open-world settings. More visualization results can be found at DetAny3D project page.
@@ 4.1 Experimental Setup
DA3D Benchmark. We present DA3D, a unified 3D detection dataset that aggregates 16 diverse datasets for 3D detection and depth estimation. Building upon Omni3D’s original datasets (Hypersim [57], ARKitScenes [4], Objectron [2], SUNRGBD [61], KITTI [24], and nuScenes [9]), we incorporate additional four outdoor detection datasets (Argoverse2 [70], A2D2 [25], Waymo [62], Cityscapes3D [21]), one indoor detection dataset (3RScan [65]), and five depth and intrinsic datasets (Scannet [17], Taskonomy [77], DrivingStereo [71], Middlebury [59], IBIMS-1 [34]). All data is standardized with monocular images, camera intrinsics, 3D bounding boxes, and depth maps. Following prior work [74], we select partial categories from KITTI, SUNRGBD, and ARKitScenes as zero-shot test classes. We select Cityscapes3D, Waymo, and 3RScan as our zero-shot datasets with novel camera configurations, where 3RScan also contains novel object categories. Depth supervision from LiDAR, RGB-D, and stereo sensors enhances 75% of training samples, while intrinsic parameters cover 20 camera configurations across 0.4 million frames (2.5× Omni3D’s scale). Dataset statistics and splits are detailed in Section 6.
@@ 6 DA3D
The dataset is standardized with the similar structure to Omni3D [8], including monocular RGB images, camera intrinsics, 3D bounding boxes, and depth maps. Omni3D++ is designed to test 3D detection models across a wide variety of environments, camera configurations, and object categories, offering a more comprehensive evaluation setting.
@@ 7.3 Loss Details
Depth Loss. The depth module is supervised using the Scale-Invariant Logarithmic (SILog) loss [], defined as:
@@ 9.1 Various Prompts Performance
In this section, we evaluate different types of prompts, including box prompts, point prompts, and text prompts, both with and without intrinsic prompts. The results on Omni3D are presented in Table 4. Each prompt type demonstrates its effectiveness in guiding 3D detection. Notably, on the zero-shot datasets, we observe that omitting intrinsic prompts leads to a significant performance drop (even approaching zero), which further highlights the critical role of intrinsic prompts for reliable depth calibration in unseen scenarios.
@@ 10 Licenses and Privacy
All the data is under the CC BY-NC-SA 4.0 license [^1] https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en. Other datasets (including nuScenes [9], Waymo [62], etc). For videos from YouTube, permission to access the video content is received through a Creative Commons license. Besides, we skip channel-related content at the beginning and end of the videos during data processing to ensure we do not infringe upon the rights of logos, channel owner information, or other copyrighted materials.

### v2

@@ Abstract
Despite the success of deep learning in close-set 3D object detection, existing approaches struggle with zero-shot generalization to novel objects and camera configurations. We introduce DetAny3D, a promptable 3D detection foundation model capable of detecting any novel object under arbitrary camera configurations using only monocular inputs. Training a foundation model for 3D detection is fundamentally constrained by the limited availability of annotated 3D data, which motivates DetAny3D to leverage the rich prior knowledge embedded in extensively pre-trained 2D foundation models to compensate for this scarcity. To effectively transfer 2D knowledge to 3D, DetAny3D incorporates two core modules: the 2D Aggregator, which aligns features from different 2D foundation models, and the 3D Interpreter with Zero-Embedding Mapping, which mitigates catastrophic forgetting in 2D-to-3D knowledge transfer. Experimental results validate the strong generalization of our DetAny3D, which not only achieves state-of-the-art performance on unseen categories and novel camera configurations, but also surpasses most competitors on in-domain data. DetAny3D sheds light on the potential of the 3D foundation model for diverse applications in real-world scenarios, e.g., rare object detection in autonomous driving, and demonstrates promise for further exploration of 3D-centric tasks in open-world settings. More visualization results can be found at our code repository.
@@ 4.1 Experimental Setup
DA3D Benchmark. We present DA3D, a unified 3D detection dataset that aggregates 16 diverse datasets for 3D detection and depth estimation. Building upon Omni3D’s original datasets (Hypersim [57], ARKitScenes [4], Objectron [2], SUNRGBD [61], KITTI [24], and nuScenes [9]), we incorporate additional four outdoor detection datasets (Argoverse2 [70], A2D2 [25], Waymo [62], Cityscapes3D [21]), one indoor detection dataset (3RScan [65]), and five depth and intrinsic datasets (Scannet [17], Taskonomy [77], DrivingStereo [71], Middlebury [59], IBIMS-1 [34]). All data is standardized with monocular images, camera intrinsics, 3D bounding boxes, and depth maps. Following prior work [74], we select partial categories from KITTI, SUNRGBD, and ARKitScenes as zero-shot test classes. We select Cityscapes3D, Waymo, and 3RScan as our zero-shot datasets with novel camera configurations, where 3RScan also contains novel object categories. Depth supervision from LiDAR, RGB-D, and stereo sensors enhances 75% of training samples, while intrinsic parameters cover 20 camera configurations across 0.4 million frames (2.5× Omni3D’s scale). Dataset statistics and splits are detailed in Section 6. All data are subject to their respective licenses.
@@ 6 DA3D
The dataset is standardized with the similar structure to Omni3D [8], including monocular RGB images, camera intrinsics, 3D bounding boxes, and depth maps. DA3D is designed to test 3D detection models across a wide variety of environments, camera configurations, and object categories, offering a more comprehensive evaluation setting.
@@ 7.3 Loss Details
Depth Loss. The depth module is supervised using the Scale-Invariant Logarithmic (SILog) loss, defined as:
@@ 9.1 Various Prompts Performance
In this section, we evaluate different types of prompts, including box prompts, point prompts, and text prompts, both with and without intrinsic prompts. The results on Omni3D are presented in Table 4. Each prompt type demonstrates its effectiveness in guiding 3D detection. Besides, on the zero-shot datasets, we observe that omitting intrinsic prompts leads to a significant performance drop (even approaching zero), which further highlights the critical role of intrinsic prompts for reliable depth calibration in unseen scenarios.
@@ 10 Licenses and Privacy
All data used in this work are obtained from publicly available datasets and are subject to their respective licenses.